Mrs Windell spent 1 h addressing one-third of the family christmas cards, and Mr. Windell spent 1 ¼ h addressing another third. Continuing the addressing together, how long did they take to finish the rest of the cards?
Answer by mananth(16949)

Mrs Windell spent 1 h addressing one-third of the family christmas cards,
Mr. Windell spent 1 ¼ h addressing another third.
..
Mr. Windell took 5/4 hours for 1/3 job
so she will do in 1 hour (1/3 )/(5/4)
=1/3*4/5
= 4/15 of the job in 1 hour.
..
together they can do 1/3 + 4/15 of the job in 1 hour
=5+4 / 15
=9/15
=3/5 of the job in 1 hour.together
..
They have already completed 1/3 +1/3 = 2/3 of the job.
so balance job is 1-2/3 =1/3 job
..
3/5 of job they take 1 hour
1/3 of job they take (1/3) / (3/5)
1/3*5/3= 5/9 hours